OpenStudy (anonymous):

prove that: tan theta +sec theta -1/tan theta - sec theta +1=1 + sin theta/cos theta

Parth (parthkohli):

Ok, never mind... I was writing something else. For reference, you can multiply both the numerator and denominator by \(\cos\theta\).

OpenStudy (anonymous):

on lhs or rhs???

Parth (parthkohli):

\[= \dfrac{\sin\theta +1 - \cos\theta}{\sin\theta - 1 + \cos\theta}\]

Parth (parthkohli):

On the LHS.

OpenStudy (anonymous):

what is cos(tan) and cos(sec)?

Parth (parthkohli):

\[= \dfrac{\left(\sin \theta + 1 - \cos\theta\right)^2}{(\sin\theta + \cos\theta - 1)(\sin\theta - (\cos\theta-1))}\]Simplify the denominator first separately\[\sin^2\theta - \left(\cos^2\theta + 1 + 2\cos\theta\right)= 1 - \cos^2\theta - \cos^2\theta - 1 - 2\cos\theta = -2\cos\theta(\cos\theta + 1)\]

Parth (parthkohli):

Remember that since tan = sin/cos, it means that tan * cos = sin.

OpenStudy (anonymous):

oh.... yes thank you!

Parth (parthkohli):

All this question involves is using the identity \(\sin^2\theta + \cos^2\theta = 1\). Just expand all the perfect squares until you get something in the form \(\sin^2 \theta + \cos^2 \theta \)

OpenStudy (anonymous):

ya ok! i'll do it! thanks for ur help!!! :)

Parth (parthkohli):

If you know the identity \( 1 + \tan^2 A = \sec^2 A\), you can use that too. But all identities are related with each other so just go ahead with any one of them.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ty!

Parth (parthkohli):

No problem :)

OpenStudy (anonymous):

hello! how do u simplify the numerator??

OpenStudy (anonymous):

plz help!!!!!!